OpenStudy (anonymous):

equation of a line perpendicular to y=3x-1?

OpenStudy (anonymous):

perpendicular line is the negative reciprocal so the slope of the line has to be \[-\frac{1}{3}\] since there's no point, you can put any constant or you can just leave it at zero \[y = -\frac{1}{3}x\]

OpenStudy (anonymous):

what do you mean since there's no point? what about the -1?

myininaya (myininaya):

hes says your question did't say what point the perpendicular line was running through

OpenStudy (anonymous):

the -1 is the intercept for y = 3x -1 but the question asks you for equation of A line, so ANY line it can be y = -(1/3)x or y = -(1/3)x + 3 y = -(1/3)x + 5 it can be any number because the question doesn't specify a specific point that the line has to go through

OpenStudy (anonymous):

it says it has to pass through the point of intersection from these two equations: y=24/x and y=3x-1

myininaya (myininaya):

y=-(1/3)x+C where C is a constant is the general perpendicular line for the line y=3x+K where K can be any constant

OpenStudy (anonymous):

it has to pass through the point where those two equations intersect?

OpenStudy (anonymous):

yes

myininaya (myininaya):

set 24/x=3x-1

myininaya (myininaya):

solve for x

OpenStudy (anonymous):

so set those equations equal to one another and solve for the point 24/x = 3x - 1 3x^2 -x = 24 3x^2 - x - 24 = 0 x = -(8/3), 3 myininaya, what am i doing wrong? how can i get two intersections for two lines???

myininaya (myininaya):

24/x is not a line its a hyperbola

OpenStudy (anonymous):

oh wait yeah, i didn't catch that..

myininaya (myininaya):

so since we have two intersections and your question didn't specify which intersection we have 2 equations: they both have the the form y=-(1/3)x+C so we have 1 equation with x=-8/3 so y=3(-8/3)-1=-9 so we have -9=-(1/3)(-8/3)+b solve for b to find the y intercept and we also have equation 2 with x=3 so y=3(3)-1=8 so we have 8=-(1/3)(3)+b solve this for b and you have your two equations

myininaya (myininaya):

so the first equation we have -9-8/9=b so b=-89/9 so equation 1 is y=-(1/3)x-89/9 and the second equation we have 8+1=b so b=9 so equation 2 is y=-(1/3)x+9

OpenStudy (anonymous):

thank you !

myininaya (myininaya):

your welcome. what class is that? is that really algebra? That is hard question for algebra students. no offense. i would know since I teach algebra

myininaya (myininaya):

i like it though. I might give it as a project hehe

OpenStudy (anonymous):

ib math studies :)

myininaya (myininaya):

ib?

OpenStudy (anonymous):

International Baccalaureate its a higher level of math offered worldwide

myininaya (myininaya):

i am jealous

myininaya (myininaya):

i wish i was challenged more but I guess I know enough lol, but no one could really know enough so I take that back

OpenStudy (anonymous):

its like college level math. well all of my courses are ib.

myininaya (myininaya):

so did you understand everything above?

OpenStudy (anonymous):

this is supposed to be an easy problem compared to the other problems we learn lol yes i did thank you !

myininaya (myininaya):

i really liked your problem. its nice